OpenStudy (gabylovesyou):

Choose the correct simplification of the expression @jim_thompson5910

OpenStudy (mertsj):

Use this property: \[\frac{1}{a ^{-n}}=a^n\]

OpenStudy (mertsj):

Or to put it in words you could say...if a factor has a negative exponent and is in the denominator, take it to the numerator and make the exponent positive.

OpenStudy (mertsj):

You need to replace both n's with the 5

OpenStudy (mertsj):

\[\frac{1}{a ^{-5}}=a^5\]

OpenStudy (mertsj):

Yes. That's it. So now your answer is: \[8a^5b\]